# Multi-Level Damage-Aware Graph Learning for Resilient UAV Swarm Networks

### v2

@@ IV-E1 Complexity of MBDA
The matrix operation in MBDA module is the Hadamard product in (14). Since the $\bm{A}_{in}^{k}$ and $\bm{M}_{da}$ are sparse, the computational complexity of the k-hop branch is determined as $\mathcal{O}\left(\text{nnz}(\bm{A}_{in}^{k})\cdot\text{nnz}(\bm{M}_{da})\right)$, where $\text{nnz}(\cdot)$ is the number of non-zero elements in a matrix. Assuming that $N$ nodes are randomly distributed within a $D\times D$ square area, the node degree follows a Poisson distribution. For the 1-hop branch, the number of non-zero elements in $\bm{A}_{in}^{1}$ can be calculated as
@@
where $p=N_{D}/N$ denotes the damage ratio. Therefore, for the 1-hop branch of the MBDA module, the entire complexity is $\mathcal{O}\left(2p(1\!-\!p)\bar{d}N^{3}\right)$. For the $k$-hop branch, the node degree still follows the Poisson distribution, hence we have $\text{nnz}(\bm{A}_{in}^{k})=k\bar{d}N$. The computational complexity of the MBDA module is finally determined as $\mathcal{O}\left(p(1\!-\!p)K(K\!-\!1)\bar{d}N^{3}\right)$.
In summary, the MBDA module requires a computational complexity of $\mathcal{O}\left(K^{2}N^{3}\right)$. Since the MBDA module runs only once for each CNS issue, it can be regarded as a data pre-processing stage, and the complexity is acceptable.
@@ IV-E2 Complexity of DGCN
The matrix operations in DGCN are mainly matrix multiplications in (24).
@@
The computational complexity of the DGCN is given by
@@
The sparsity of the B-GCO kernel provides the DGCN a linear complexity of $\mathcal{O}\left(K^{2}N\right)$ according to Proposition 2. Therefore, the proposed algorithm has computational complexity of $\mathcal{O}\left(K^{2}N^{3}+RK^{2}N\right)$, where $R$ is the round of online iterations. For actual practice, the proposed algorithm needs 6.02 Mb parameters and 513 MiB running memory for USNET with 200 UAVs, showing acceptable complexity for practical deployment.
@@ V Simulation Results
This section first introduces the setup of simulation experiments  [^1] The source codes of the proposed ML-DAGL algorithm are available on https://github.com/lytxzt/Damage-Attentive-Graph-Learning.. Then, comparative results with other advanced approaches are presented in the second subsection, and ablation studies are discussed in the third subsection.
@@ V-A Simulation Setup
V-A1 Inferences
The original USNET, consisting of $N$ UAVs, was randomly distributed within a $D\times D$ square area, where the node density is set to $\rho=200\,\text{node}/km^{2}$. Based on the analysis in [23], we set the communication distance as $d_{tr}=120\,m$. The maximum value of UAV velocity was defined as $v_{max}=10\,m/s$ for ease of calculation, and each recovery step had a time interval $\Delta t=0.1{\rm s/step}$. The maximum recovery time is set to $T_{rc}^{max}={D}/{2v_{max}}$. The number of destroyed UAVs is given by $N_{D}=pN$ according to the damage ratio $p$. For the hyperparameter in the proposed algorithm, the number of residual blocks in the DGCN is $L=3$, and the hidden dimension is $d_{s}=512$. The learning rate is set to 0.0001 with an Adam optimizer, and the dropout ratio is set to 0.1 to alleviate overfitting. The pre-trained model is determined as the parameters after 500 iterations under a random damage case for each $N$ with $p=0.5$.
V-A2 Metrics
Three metrics for performance evaluation were presented in this paper. The convergent ratio $R_{c}$ represents the proportion of the number of experiments that complete connectivity recovery within $T_{rc}^{max}$ to the total number of experiments. The higher $R_{c}$ indicates a more robust adaptability for different scenarios. The recovery time $T_{rc}$ is the average time for finishing the recovery process to evaluate the performance of different resilient algorithms. Moreover, we introduce degree distribution $P_{d}$ to measure the topology uniformity of the final recovered $\mathcal{\hat{G}}_{R}$, which denotes the proportion of remaining nodes with a degree number not greater than $d$ versus the total number of nodes in $\mathcal{\hat{G}}_{R}$. A more uniform degree distribution leads to a network that is more robust against attacks [31].
V-B Comparative Results Evaluation
To comprehensively evaluate the performance of our proposed algorithm, we compare it with several advanced resilient approaches, including the traditional algorithms (direct-centering, HERO [21], SIDR [22]) and the graph-learning algorithms (original GCN [25], GAT[32], CR-MGC [23], DEMD [24]).
V-B1 Scalability on swarm scales $N$
To evaluate the scalability of the proposed algorithm for varying swarm scales, we set the number of UAVs as $N\in\{50,100,200,500,1000\}$. The damage ratio is set to $p=0.5$, and each simulation of different $N$ is repeated 50 times randomly for the statistical results.
TABLE I: Statistical results of algorithms under half-damaged scenarios with different swarm scales.
  Algorithm $N=20$ $R_{c}\,/\,E(T_{rc})\,/\,\sigma(T_{rc})\,/\,E(d)\,/\,max(d)$ $N=50$ $R_{c}\,/\,E(T_{rc})\,/\,\sigma(T_{rc})\,/\,E(d)\,/\,max(d)$ $N=100$ $R_{c}\,/\,E(T_{rc})\,/\,\sigma(T_{rc})\,/\,E(d)\,/\,max(d)$
[…]
  Algorithm $N=200$ $R_{c}\,/\,E(T_{rc})\,/\,\sigma(T_{rc})\,/\,E(d)\,/\,max(d)$ $N=500$ $R_{c}\,/\,E(T_{rc})\,/\,\sigma(T_{rc})\,/\,E(d)\,/\,max(d)$ $N=1000$ $R_{c}\,/\,E(T_{rc})\,/\,\sigma(T_{rc})\,/\,E(d)\,/\,max(d)$
centering 0.98/ 22.90/ 11.17/ 26.54/ 980.9822.9011.1726.54980.98\,/\,22.90\,/\,11.17\,/\,26.54\,/\,980.98 / 22.90 / 11.17 / 26.54 / 98 0.88/ 46.70/ 18.81/ 71.04/ 2380.8846.7018.8171.042380.88\,/\,46.70\,/\,18.81\,/\,71.04\,/\,2380.88 / 46.70 / 18.81 / 71.04 / 238 0.50/ 90.12/ 26.16/ 260.78/ 4630.5090.1226.16260.784630.50\,/\,90.12\,/\,26.16\,/\,260.78\,/\,4630.50 / 90.12 / 26.16 / 260.78 / 463
HERO 0////0\,/\,-\,/\,-\,/\,-\,/\,-0 / - / - / - / - 0////0\,/\,-\,/\,-\,/\,-\,/\,-0 / - / - / - / - 0////0\,/\,-\,/\,-\,/\,-\,/\,-0 / - / - / - / -
[…]
GCN 0.86/ 30.12/ 12.72/ 12.14/ 630.8630.1212.7212.14630.86\,/\,30.12\,/\,12.72\,/\,12.14\,/\,630.86 / 30.12 / 12.72 / 12.14 / 63 0.86/ 58.52/ 14.52/ 56.72/ 2140.8658.5214.5256.722140.86\,/\,58.52\,/\,14.52\,/\,56.72\,/\,2140.86 / 58.52 / 14.52 / 56.72 / 214 0.08////0.08\,/\,-\,/\,-\,/\,-\,/\,-0.08 / - / - / - / -
[…]
Simulation results of the convergent ratio $R_{c}$, mean recovery time $E(T_{rc})$, variance of recovery time $\sigma(T_{rc})$, average node degree $E(d)$, and maximum node degree $max(d)$ with different resilient algorithms are presented in TABLE I. For the convergence $R_{c}$ of each resilient algorithm, it can be observed that the proposed ML-DAGL can always guarantee the restoration of connectivity under varying scales, while other algorithms have decreasing $R_{c}$ and lack the scalability on large swarms, especially when $N$ reaches 1000. For the recovery time, ML-DAGL achieves state-of-the-art performance on both means and variances of $T_{rc}$, showing an average of 54.9% reduction compared to previous algorithms, and more stable solutions for CNS issues under different $N$. In terms of the node degree distribution, SIDR has the lowest $E(d)$ and $max(d)$ owing to its distributed recovery implementation. Over the graph learning-based recovery algorithms, ML-DAGL achieves a lower-degree distribution. This indicates that the proposed algorithm solves the challenge of over-aggregation, significantly improving the topologies of recovered USNETs.
V-B2 Adaptability on damage scales $N_{D}$
To further explore the adaptability of the proposed algorithm under varying damage scales, we randomly destroy 10 to 190 UAVs with the original USNET set as $N=200$, distributed within a $1\,km\times 1\,km$ area. The maximum recovery time is set to $T_{rc}^{max}=50\,s$, and similarly, each destruction experiment is repeated 50 times for statistical results.
The convergent ratios of each resilient algorithm are illustrated in Fig.4. The convergence of HERO, SIDR, GCN, and GAT decreases rapidly as the number of destroyed UAVs increases, while the convergent ratios of centering, CR-MGC, and DEMD, remain close to 1 under varying damage scenarios. However, these algorithms cannot guarantee the connectivity when $N_{D}$ exceeds 150. For the proposed algorithm, the convergent ratio always equals 1, which demonstrates that all cases of CNS issues are solved. Therefore, the proposed algorithm has better convergence on sparse topologies caused by extremely massive damage.
The average recovery time $T_{rc}$ of different resilient algorithms under damage of different $N_{D}$ are shown in Fig.5, where the shaded areas represent the 99% confidence intervals of the average recovery time. We can see that larger damage scales $N_{D}$ require longer recovery time across all approaches. Once the damage exceeds 100 nodes, this growth trend accelerates significantly. Referring to the comparative results, the proposed ML-DAGL algorithm shows notable performance improvements in recovery time with an average of 59.8% reduction compared with existing approaches. Specifically, for the half-damaged scenario with $N_{D}=100$, the average recovery time of the proposed algorithm is cut by 75.4% compared with CR-MGC and DEMD, which is a great margin of 16 seconds. Therefore, the proposed algorithm can accelerate the recovery process of CNS issues under varying damage scenarios.
Figure 4: Results of convergent ratio $R_{c}$ under different numbers of destroyed UAVs $N_{D}$.
Figure 5: Results of average recovery time $T_{rc}$ under different numbers of destroyed UAVs $N_{D}$.
@@ V-B2 Adaptability on damage scales $N_{D}$
Figure 6: Results of cumulative degree distribution $P_{d}$ under different numbers of destroyed UAVs $N_{D}$.
The cumulative node-degree distributions under 100 and 150 destroyed UAVs with $N=200$ are illustrated in Fig.6a and Fig.6b. The results indicate that the degree distributions of the proposed algorithm are primarily concentrated in low-degree regions, with a faster-increasing curve. In contrast, other approaches tend to have higher-degree nodes within the recovered topologies. Therefore, the ML-DAGL helps to rebuild a network with a more sparse and uniform topology after recovery, and the absence of large-degree nodes also improves the swarm resilience when the USNET faces the risk of secondary damage.
V-B3 Case study
@@
To compare the performance of different resilient algorithms more intuitively, we introduce a specific example to demonstrate the network recovery process based on each approach. As shown in Fig.7a, we randomly destroy $N_{D}=100$ UAVs of the original USNET, and the remaining graph is divided into 7 disconnected sub-nets with different sizes, leading to an uneven topology. The numbers of sub-nets during the recovery process by different algorithms are illustrated in Fig.7b. It takes only 2.4s for the proposed algorithm to rebuild a connected remaining graph, while other algorithms require much more recovery time. Moreover, the recovery trajectories of CR-MGC, DEMD, and ML-DAGL are plotted respectively. Obviously, the remaining nodes in Fig.7c and Fig.7d tend to aggregate towards sub-net 1 as it is the largest sub-net with the highest degrees, thus causing the over-aggregation problem. The trajectories of ML-DAGL in Fig.7e are quite shorter, and the moving directions of the remaining nodes are towards the damaged regions nearby. Therefore, the proposed algorithm can provide solutions with wider spatial distributions and more uniform topologies, thereby improving communication coverage and alleviating the problem of local over-aggregation.
V-B4 Response time of algorithms
Figure 8: Results of average time consumption of different algorithms under $N=200$.
The average time consumption of different algorithms is compared in Fig.8. It is worth noting that the centering, HERO, and SIDR require online calculation on every execution step, hence, the entire time consumption of these algorithms is related to their time cost of the recovery process. The graph-learning algorithms only need to calculate the recovery solution when the damage occurs. Compared with other graph-learning approaches, the proposed algorithm has acceptable time consumption for addressing CNS issues. This is because the ML-DAGL requires fewer iterations of GCO thanks to the pre-trained model, which is discussed in a later subsection of the ablation study.
V-C Ablation Study
@@
V-C1 Study of mDAG sequence
@@
We first consider the damage-attention in the proposed algorithm. Although the comparative results proved that mDAG can accelerate the recovery process and mitigate the over-aggregation problem, there is still a great advantage of utilizing mDAG sequence as the input of GCN, that is, it uses a unified input matrix framework. Therefore, when the number of UAVs in the original USNET is fixed, only one pre-trained model is necessitated which can adapt to damage scenarios with different $N_{D}$. This significantly reduces storage needs for pre-trained parameters and decreases the overhead of online iterative computing. The loss curves of ML-DAGL during the training process are plotted in Fig.9, where the weights $W$ in (24) are initialized by pre-trained model and random parameters, respectively. The loss curves initialized by the pre-trained model started from smaller loss values and achieved a smooth stage at the 50th episode. This means that using a pre-trained model can reduce the time consumption for addressing CNS issues. Meanwhile, the same pre-trained model can initialize the ML-DAGL under different $N_{D}$ with similar curve trends. Therefore, the proposed algorithm has less storage requirement of only one pre-trained model, while CR-MGC needs 199 meta-learning models.
V-C2 Study of MBDA
The proposed MBDA module offers two advantages to ML-DAGL. First, it applies multi-hop dilation to extend the receptive fields of nodes under massive damage scenarios. Second, it has a structure of parallel dilation branches, which allows the parallel computation of solutions $\bm{\hat{X}}_{gcn}^{k}$ with different dilation sizes $k$, optimizing the recovery path for varying damage scales. The recovery time of the MBDA with $k$-hop and parallel dilation is shown in Fig.10. The results demonstrate that the optimal value of dilation size $k$ corresponds to the scale of destruction, e.g., the best choice of $k$ is $k=3$ when $N_{D}=100$ while it becomes $k=9$ when $N_{D}=150$. Compared with specific dilation size $k$, the parallel dilation has a shorter recovery time across all $N_{D}$. Although the parallel structure increases the number of parameters, the ML-DAGL algorithm has only a 6.02Mb model size that is manageable given the performance gains achieved.
@@ V-C2 Study of MBDA
V-C3 Study of DGCN
The proposed DGCN has two key components for generating solutions to CNS issues, i.e., the B-GCO and the residual connection mechanism. The average recovery time of ablation experiments on DGCN is plotted in Fig.11, where the B-GCO is replaced by the original GCO in [25] and the residual connection is optional. The results indicate that DGCN without the residual connection requires a longer recovery time when $N_{D}$ gets larger, as the convolution becomes inefficient on a sparse topology. This positively shows the effectiveness of the residual connection under extremely massive damage. Meanwhile, the B-GCO is outperforming the original GCO with an average 22.4% margin. This is because the original GCO normalizes the degree at the node level, while B-GCO uses degree normalization at the graph level, which allows nodes to retain the original degree weight information. As a result, the DGCN can efficiently utilize the original network information to address CNS issues.

### v3

@@ IV-E1 Complexity of MBDA
Matrix operations in MBDA module is the Hadamard product in (15). Although MBDA enhance the sparsity of $\bm{A}_{in}^{k}$, the damage attention matrix $\bm{M}_{da}$ is always a sparse matrix, hence the computational complexity of the k-hop branch is calculated as $\mathcal{O}\left(\text{nnz}(\bm{A}_{in}^{k})\cdot\text{nnz}(\bm{M}_{da})\right)$, where $\text{nnz}(\cdot)$ is the number of non-zero elements in a matrix.
For the 1-hop branch, assuming that $N$ nodes are randomly distributed within a $D\times D$ square area, two nodes have a probability of $\frac{\pi d_{tr}^{2}}{D^{2}}$ to become neighbors, hence the node degree follows a Poisson distribution. The number of non-zero elements in $\bm{A}_{in}^{1}$ can be calculated as
@@
where $p=N_{D}/N$ denotes the damage ratio. Therefore, for the 1-hop branch of the MBDA module, the entire complexity is $\mathcal{O}\left(2p(1\!-\!p)\bar{d}N^{3}\right)$.
For the $k$-hop branch, we can regard the $k$-hop dilation as a node extends its communication range $k$ times, hence the node degree in $\bm{A}_{in}^{k}$ still follows the Poisson distribution, hence we have $\text{nnz}(\bm{A}_{in}^{k})=k^{2}\bar{d}N$. The computational complexity of the MBDA module is finally determined as $\mathcal{O}\left(\frac{1}{3}p(1\!-\!p)K(K\!-\!1)(2K\!-\!1)\bar{d}N^{3}\right)$.
Consequently, the MBDA module requires a computational complexity of $\mathcal{O}\left(K^{3}N^{3}\right)$. Since the MBDA module is a data pre-processing stage, it runs only once for each CNS issue with necessary complexity.
@@ IV-E2 Complexity of DGCN
For DGCN, the matrix operations are mainly matrix multiplications in B-GCO layers. The total complexity is related to the number of B-GCO layers and the round of iterations. Therefore, we first determine the complexity of DGCN within one iteration.
@@
The one-round computational complexity of the DGCN is given by
@@
Therefore, the DGCN has a linear computational complexity of $\mathcal{O}\left(K^{2}N\right)$ to $N$ for a single iteration. This is because the sparsity of the B-GCO kernel provides a significant reduction in the complexity of matrix multiplication during graph convolutions, according to Proposition 2.
In summary, the proposed algorithm has total computational complexity of $\mathcal{O}\left(K^{3}N^{3}+RK^{2}N\right)$, where $R$ is the round of online iterations. With the pre-trained model, ML-DAGL algorithm requires a small $R$ to compute the optimal solution, hence the computational complexity is acceptable.
@@ V Simulation Results
This section first introduces the setup of simulation experiments and then presents the performance of the proposed algorithm  [^1] The source codes of the proposed ML-DAGL algorithm are available on https://github.com/lytxzt/Damage-Attentive-Graph-Learning. The computations in this paper were run on the $\pi$ 2.0 cluster supported by the Center for High Performance Computing at Shanghai Jiao Tong University.. The performances of center-fly, HERO [21], SIDR [22]), and the graph-learning algorithms including original GCN [25], GAT[32], CR-MGC [23], DEMD [24], are also displayed for comparisons.
@@ V-A Simulation Setup
V-A1 Environment initialization
In the simulation, the original USNET is composed of $N$ UAVs randomly deployed within a square area of size $D\times D$, with a fixed node density of $\rho=200\,\text{nodes/km}^{2}$. A fraction of the UAVs is assumed to be destroyed due to attacks or failures, where the number of failed UAVs is given by $N_{D}=pN$, with $p$ representing the damage ratio. The maximum allowed recovery time is set to be the time cost of flying across the half area, i.e., $T_{rc}^{max}=D/(2v_{max})$. Solutions fail to restore connectivity within $T_{rc}^{max}$ are regard as not convergent.
For constant parameters, the communication radius is set to $d_{tr}=120\,\text{m}$ following the analysis in [23]. The maximum UAV velocity is constrained to $v_{max}=10\,\text{m/s}$ for computational convenience. Recovery is modeled in discrete time steps with a fixed interval of $\Delta t=0.1\,\text{s/step}$.
V-A2 Model Parameters
We develop $L=3$ residual blocks in DGCN and set the hidden feature dimension to $d_{s}=512$. The Adam optimizer with a learning rate of 0.0001 is applied for training and the dropout rate is set to 0.1 to prevent overfitting. The pre-trained models are trained for 500 iterations under randomly $p=0.5$ damage scenarios with different swarm scale $N=20,50,100,200,500,1000$.
V-A3 Evaluation Metrics
Three metrics are adopted to evaluate the performance of resilience. The convergent Ratio $R_{c}$ denotes the proportion of simulations that can restore connectivity within the allowed time $T_{rc}^{max}$, which reflects the robustness of algorithms. The recovery Time $T_{rc}$ is the average time required to complete the connectivity restoration process, offering a direct measure of the algorithm’s efficiency. The degree Distribution $P_{d}$ denotes the proportion of nodes whose degrees do not exceed $d$ in the recovered network. The mean degree $E(d)$ and maximum degree $max(d)$ also represent the distribution characteristics.
@@ V-A3 Evaluation Metrics
((c))
((d))
Figure 4: Statistical results of resilient algorithms under varying damage scenarios with swarm scale $N=200$.
V-B Statistical Results Evaluation
We develop $N=200$ UAVs within a $1\,\text{km}\times 1\,\text{km}$ area and randomly destroy between 10 and 190 UAVs to evaluate the performance of the proposed algorithm under varying damage levels. Therefore, The period of recovery process is $T_{rc}^{max}=50\,\text{s}$. Each scenario is repeated 50 times for statistical robustness. Simulation results for resilient algorithms are presented in Fig.4.
The convergent ratios $R_{c}$ of resilient algorithms are illustrated in Fig.4(a), showing that $R_{c}$ for HERO, SIDR, GCN, and GAT declines sharply with increasing $N_{D}$, while center-fly, CR-MGC, and DEMD maintain near-perfect convergence up to $N_{D}=150$. In contrast, the proposed ML-DAGL consistently achieves $R_{c}=1$, successfully resolving all connectivity restoration cases, including under extreme damage scenarios.
The average recovery time $T_{rc}$ verses $N_{D}$ are plotted in Fig.4(b). It can be observed that all approaches require longer $T_{rc}$ when $N_{D}$ increases. The curve of ML-DAGL algorithms first grows slowly then increase rapidly after $N_{D}$ reaching 100. On the other hand, ML-DAGL significantly outperforms other approaches, reducing average recovery time by 59.8% compared with second-best approach. For instance, the proposed algorithm achieves 81.7%, 71.0%, and 30.0% reductions compared to DEMD under $N_{D}=50$, 100, and 150, respectively. These results confirm ML-DAGL’s efficiency across varying damage scales.
The cumulative node-degree distributions $P(d)$ for $N_{D}=100$ and $150$ are shown in Fig.4(c) and Fig.4(d). The degree of recovered network by ML-DAGL is distributed in lower region, while other approaches have more high-degree nodes. This demonstrates that the proposed algorithm mitigates the over-aggregation and thereby forms more even topologies.
V-C Case Study
@@
To provide a more intuitive comparison of the recovery performance across different resilient algorithms, we present an example illustrating the network restoration process under each approach as the case study, as shown in Fig.5.
In Fig.5(a), $N_{D}=100$ UAVs are randomly removed from the original USNET, resulting in a split topology with 7 disconnected sub-nets of different sizes. The numbers of sub-nets during recovery are plotted in Fig.5(b). The proposed ML-DAGL algorithm restores connectivity in just 2.4 seconds, whereas other approaches require substantially more time.
The recovery trajectories produced by CR-MGC, DEMD, and ML-DAGL are shown in Fig.5(c) to Fig.5(e), respectively. We can see that the trajectories of CR-MGC and DEMD tend to converge toward sub-net 1, which has the largest size and highest degrees for aggregation. In contrast, the trajectories under ML-DAGL are notably shorter, with UAVs moving toward nearby damaged regions rather than dominant clusters. This leads to more distributed solutions and uniform topologies, thereby mitigating local over-aggregation.
V-D Online Iteration Consumption
Figure 6: Results of total online average time consumption of different algorithms under $N=200$.
The average time consumptions of online iteration with different algorithms are presented in Fig.6. Note that center-fly, HERO, and SIDR perform online computations at each execution step, their total time consumptions get pretty high under pro-longed recovery processes. In contrast, graph learning-based algorithms compute the recovery strategy only once upon, and the proposed ML-DAGL demonstrates competitive time cost for solving CNS issues. For practice, ML-DAGL needs only 6.02 Mb parameters and 513 MiB running memory for USNET with 200 UAVs, showing acceptable complexity for practical deployment.
V-E Ablation Study
@@
V-E1 Study of pre-trained model
@@
We represent the loss curves of ML-DAGL during training in Fig.7, where the weights $W$ in (25) are initialized using either a pre-trained model or random parameters. The results show that the curves with the pre-trained weights begin with lower loss values and reach a stable phase at the 50th episode. In contrast, curves of random initialization require more iterations to converge according to $N_{D}$. Furthermore, we can see that the same pre-trained model works regardless of variations in $N_{D}$, indicating that only a single pre-trained model for a given $N$ is necessary. The implementation of the pre-trained model can significantly reduce the training time with limited storage requirements.
V-E2 Study of MBDA
The proposed MBDA module utilizes $K$ branches for receptive field dilation. We illustrate the recovery results of MBDA with a single $k$-th branch or parallel branches in Fig.8. The results indicate that the optimal branches $k$ is dependent on the extent of damage. For example, $k=3$ yields the best performance when $N_{D}=100$, while $k=9$ is optimal for $N_{D}=150$. Compared to any fixed dilation size $k$, the parallel dilation structure achieves shorter recovery times across all values of $N_{D}$. Although this parallel design introduces additional computations, the value of $K$ is limited by (36), with acceptable cost considering the substantial performance improvements it provides.
@@ V-E2 Study of MBDA
V-E3 Study of DGCN